If the price of a pizza rises from 250 to 300, and quantity demanded falls from 10,000 to
7,400, using the formula for arc price elasticity (5 marks)
a. What is the percentage change in price
b. what is the percentage change in quantity
c. What is the price elasticity of demand?
d. What will be the interpretation of the result
e. What are the determinant factors for the price elasticity of demand for the pizza?
1
Expert's answer
2021-08-20T18:04:01-0400

the formula for arc price elasticity:

"Edp=\frac{\frac{Q2-Q1}{\frac{Q1+Q2}{2}}}{\frac{P2-P1}{\frac{P1+P2}{2}}}=\frac{\Delta Q}{\Delta P}"

a."\Delta P=\frac{P2-P1}{\frac{P1+P2}{2}}=\frac{300-250}{\frac{250+300}{2}}=0.1818"


"\frac{300}{250}=1.2" or 20%


b."\Delta Q=\frac{Q2-Q1}{\frac{Q1+Q2}{2}}=\frac{7400-10000}{\frac{7400+10000}{2}}=-0.2989"

"\frac{7400}{10000}=0.74" or 26%


c."Edp=\frac{\frac{Q2-Q1}{\frac{Q1+Q2}{2}}}{\frac{P2-P1}{\frac{P1+P2}{2}}}=\frac{\Delta Q}{\Delta P}=\frac{-0.2989}{0.1818}=-1.64"


d. For all goods, with the exception of Giffen's goods, the coefficient of elasticity of demand for price is negative. Elastic demand. Buyers are sensitive to price changes. The demand changes more as a percentage than the price, that is, for elastic demand | Ed | > 1.0.


e. The price elasticity of demand shows by how many percent the value of demand will change when the price changes by 1 %. The price elasticity of demand is influenced by the following factors: the presence of competing goods or substitute goods, a noticeable change in the price level for the buyer; modernity and awareness of the buyers about the market of goods of interest; the time factor; the specific weight of the goods in the consumer's expenses